OpenStudy (anonymous):

solve the system: y=2x + 1 2X^2 + y^2 = 1

OpenStudy (anonymous):

x = -2/3, y = -1/3 x = 0, y = 1 Would you like an explanation?

OpenStudy (anonymous):

yes please!